# Inattention to States and Characteristics

### v1

@@ Abstract
We study a generalization of the state-action rational inattention model with two dimensions of uncertainty: states and hedonic characteristics. The resulting conditional choice probability takes a familiar weighted multinomial logit form, but in contrast to the state-action model, is unique and is not a corner solution. The model imposes testable restrictions on the behavior of choice probabilities in markets with product entry. The optimal information policy can be found by optimizing subject to a marginal constraint, then finding the optimal marginal. The solution to the former is the Schrödinger bridge of an entropic optimal transport problem. The latter is solved by maximizing a strictly concave function (over a convex set) which envelops the objective.
@@ Introduction
Consumer choice typically contains two dimensions of uncertainty: one concerns the product, and the second concerns utility-relevant future states of the world. When buying insurance, for example, the consumer faces uncertainty surrounding the contents of each insurance contract and their own future risk.
An increasingly popular framework for studying decision-making under uncertainty is the state-action rational inattention model [4]. In it, the decision-maker (DM) commits to an information policy consisting of a joint distribution $P$ over actions and states, subject to a Bayes plausibility constraint. The DM seeks to maximize expected utility less an information cost which is linear in the Shannon information between state and action. A relationship between the optimal conditional choice probability and the optimal marginal probability is known; however, the optimal marginal is not necessarily unique, and therefore neither is the optimal conditional.
We present a natural extension of the state-action model, which we call the state-characteristic model. In this model, the DM faces an exogenous prior distribution $\phi$ over characteristics $\xi$ – reflecting the inherent ubiquity of certain characteristics – and commits to a joint distribution $P$ over characteristics $\xi$ and states $\theta$. Mathematically, our model differs in only one respect: in addition to the usual Shannon information cost, we add a cost of diverging from $\phi$, interpretable as information gain from $P$ over $\phi$. This models, for instance, the fact that in a market full of lemons, one would need to acquire information to ensure that the likelihood of buying lemons is low.
The dual information costs account for the two dimensions of information gain that is observed by the econometrician when they observe the joint distribution over $(\xi,\theta)$. First, the econometrician observes there is information gain from $P(\xi|\theta)$ over $\phi(\xi)$ – that is, the choice distribution over characteristics is not what it would be if consumers chose at random. Second, the econometrician observes that the consumer’s choice of coverage level provides information about what their future losses will be – which differs from the prior $\mu$.
In section˜2, we motivate the need for the state-characteristic model by showing that the state-action model cannot account for both dimensions of information observed by the econometrician: modeling characteristics as actions fails to account for information gain from $P(\xi|\theta)$ over $\phi(\xi)$, while modeling characteristics as stochastic fails to account for information gain from $P(\theta|\xi)$ over $\mu(\theta)$. The state-characteristic model can be thought of as an interpolation between these two approaches which includes both information costs. We present the model in full generality in section˜3.
In our model, the relative costs for each dimension of information is parametrized by $\alpha\in[0,1]$. Our model thus nests the two aforementioned versions of the state-action model discussed in section˜2 at $\alpha=0$ and $\alpha=1$. As results have already been established for these cases, we focus on the case $\alpha\in(0,1)$.
Our main result, presented in section˜4, is very similar to the famous weighted multinomial logit (MNL) result [4]. Whereas their optimal conditional choice probability (CCP) satisfied
$P(a|\theta)=\frac{P(a)e^{u(a,\theta)/\lambda}}{Z(\theta;P)}$
for some normalizing partition function $Z$, our CCP satisfies
$P(\xi|\theta)=\frac{\phi(\xi)^{\alpha}P(\xi)^{1-\alpha}e^{u(a,\theta)/\lambda}}{Z(\theta;P)}$
One major advantage that our model ($\alpha\in(0,1)$) has is it always admits a unique solution whereas state-action ($\alpha=0$) does not. Therefore, even in cases where state-action seems to better model the situation, the state-characteristic model may nevertheless be useful for selecting from multiple solutions. Another distinction is that in state-action model, the DM forms (possibly non-unique) optimal “consideration sets” [1] – such that actions outside the set have $P(a)=0$ – whereas a key feature of the state-characteristic model is that no characteristic is completely ruled out.
This reflects the idea that consumers typically content themselves with being pretty sure of certain things; being truly 100% sure of anything requires an irrational amount of effort. For example, to acquire information as to how sausages are made, the conscious consumer may read the nutrition facts, the concerned consumer might read the ingredients, but nobody burglarizes a sausage factory with a spectrometer to make sure that the sausages are not produced in an environment with trace amounts of soy.
In lay terms, what the MNL formula states is that the CCP depends on three things. The first is the utility. The second is the inherent commonality of characteristics $\phi$: the more lemons there are, the more likely it is that you will buy one. The third is the marginal probability $P$. If you prefer $\xi$ to $\xi^{\prime}$ in 99% of states, then even in the 1% of states where you slightly prefer $\xi^{\prime}$ to $\xi$, you may still end up picking $\xi$ unless you acquire enough information about the state to be quite sure you are in that 1%.
How much each of these three things matters depends on the parameters $\alpha,\lambda$. If $\alpha\lambda$ is large, that means it is harder to learn about the characteristics $\xi$. Thus, $P(\xi|\theta)$ will resemble what would happen if you were to choose randomly, which is represented by the prior $\phi(\xi)$. If $(1-\alpha)\lambda$ is large, that means it is harder to learn about states $\theta$. Thus, $P(\xi|\theta)$ will not differ much from $P(\xi)$, since separating $P(\xi|\theta)$ from $P(\xi)$ means you learned to differentiate $\theta$ from other states. If both are large, then $P(\xi)$ is close to $\phi(\xi)$ and the CCPs are close to both. Conversely, as $\lambda\to 0$, the CCP begins to look like that of a fully-informed, utility (soft-)maximizing agent.
Lastly, section˜5 concerns falsifiable restrictions implied by the results. In particular, entry of products into the market which generate variation in $\phi$ should generate changes in CCP ratios that are not state-dependent. These predictions do not require knowledge of the prior $\phi$ to be tested. However, if the prior is known, the relative cost parameter $\alpha$ is over-identified in the process.
In the finite discrete case, the results may be established by taking Lagrangians. In the remaining sections, we derive the results for the general case, since characteristics may be continuous. The results hereafter are mostly mathematical in nature, whose utility lies in that they provide (hopefully) interesting and (hopefully) novel insights into the problem, as well as connections to other well-known problems. The proofs in these sections are typically deferred to the appendix.
Section˜6 explores the first-step orthogonality property of Shannon entropy-based inattention objectives and its consequences. In brief, the objective (which is a function of the information policy $P$) can be written as an integral in which $P$ appears twice: once in the integrand and once in the form of $dP$. While the product rule suggests that perturbing $P$ ought to affect the objective in two ways, first-step orthogonality establishes that it only locally affects the objective in one: via the re-weighting in the form of $dP$. First-step orthogonality implies that both state-action and state-characteristic have a nice property whereby when the objective is written as a single integral, the integrand is constant.
Section˜7 establishes a connection between rational inattention and entropic optimal transport (EOT). It turns out that the problem of maximizing the state-characteristic objective subject to a second marginal constraint $P|_{\xi}=\nu$ (on top of Bayes plausibility, which is itself a marginal constraint $P|_{\theta}=\mu$) can be re-cast as a Schrödinger bridge problem [7]. The problem comes with a dual which consists of finding Schrödinger potentials $a(\xi),b(\theta)$ which satisfy the associated Schrödinger equations. From EOT theory, we know that the value as a function of the marginal constraint $\nu$ is additively separable and equal to the sum of (the expectations of) the potentials. The potentials have a straightforward Lagrangian interpretation. Loosely speaking, the potential associated with the characteristic $\xi$ corresponds to how much the value would change if we were to change the marginal constraint $\nu$ such that some extra mass were added to $\xi$.
Perhaps the most interesting practical takeaway here is that the optimal CCP subject to two marginal constraints is also a weighted MNL. [4] had established that
$\text{$P$ is optimal}\implies P(\xi|\theta)\propto P(\xi)e^{u(\xi,\theta)/\lambda}$
but this does not necessarily imply that
$\text{$P$ is optimal subject to }P|_{\xi}=\nu\implies P(\xi|\theta)\propto\nu(\xi)e^{u(\xi,\theta)/\lambda}$
Indeed, it is not true. However, the Schrödinger equations imply that $P$ maximizes the objective subject to $P|_{\xi}=\nu,P|_{\theta}=\mu$ if and only if there exists $a,b$ such that
$P(\xi|\theta)=\phi(\xi)^{\alpha}\nu(\xi)^{1-\alpha}e^{[u(\xi,\theta)-a(\xi)-b(\theta)]/\lambda}$
Thus, $P(\xi|\theta)$ always follows a weighted MNL. Finding such an $a,b$ when $\xi,\theta$ are discrete is a case of the matrix scaling problem. Sinkhorn’s algorithm can be used to find the Schrödinger potentials.
We finish by considering the problem of finding the optimal marginal in section˜8. In [1], the authors write out a concave function over the set of marginals which helps them establish that the Kuhn-Tucker conditions are sufficient, as opposed to only necessary. This function may be obtained by swapping an expectation with a log; by Jensen’s inequality, it upper bounds the objective. It turns out that this function has even nicer properties when $\alpha\in(0,1)$. In particular, it is strictly concave over a convex set and coincides with the objective at their mutual optima. Thus, rather than maximize a more complicated objective, it suffices to solve this much easier convex optimization problem.
@@
State-action presents two options for modeling consumer choice with uncertain characteristics: either treat the characteristics as actions – i.e. costlessly observed – or as a state variable, alongside other state variables. The latter option, however, presents an identification problem. It turns out that a DM who does not observe the other state variables can achieve the same payoff at the same information cost as someone who costlessly observes them. Thus, the two behave identically and are observationally equivalent.
Example. Consumers choose between two insurance contracts $j=1,2$. They know one is high-coverage $\xi_{j}=H$ and the other is low-coverage $\xi_{j}=L$ but do not know which is which. The consumer has also two future health expenditure states, $t\in\{H,L\}$, each with probability $1/2$. The consumer receives utility 1 if they choose the ‘correct’ contract (high coverage iff high future expenditures) and 0 otherwise. In other words,
$u(j,\xi_{1},\xi_{2},t)=\textbf{1}\{\xi_{j}=t\}=\begin{cases}1&\xi_{j}=t\\ 0&\xi_{j}\neq t\end{cases}$
Recall the state-action sequence proceeds as follows: the consumer commits to an information policy – a Bayes plausible joint distribution over actions $j$ and states $(\xi_{1},\xi_{2},t)$. Nature then draws an action recommendation and a state. Now, suppose consumer 1 observes $t$ before committing to the information policy
if $\xi_{1}=t$, recommend $j=1$ with probability $p$ and $j=2$ with probability 1p1-p
if $\xi_{2}=t$, recommend $j=2$ with probability $p$ and $j=1$ with probability 1p1-p
and consumer 2 does not observe $t$, but commits to the same information policy. Consider:
  1. 1.
    Do the two consumers have the same information upon receiving the same recommendation $j$?
  2. 2.
    Do the two consumers pay the same information cost?
  3. 3.
    Is the expected payoff (averaged over $t$) the same?
The answer to the first, clearly, is no. Consumer 1 knows $t$ and $j$, whereas consumer 2 only gets $j$. Consumer 1 knows, with probability $p$, that the coverage level will be $t$, whereas consumer 2 merely knows that with probability $p$, the coverage level will be appropriate for their future health expenditures – but does not know what either are. The answer to the second is yes. The answer to the third is also yes – for both, their utility will be 1 with probability $p$ and 0 with 1p1-p.
Ergo, consumer 2 is able to achieve the same expected payoff at the same information cost, despite having less information before committing to the information policy and after receiving the recommendation. Worse, for the econometrician, the conditional choice probability for both consumers is the same, so they are indistinguishable.
What is happening is that both consumers are acquiring the same amount of information about whether $j$ is the “right” choice. Consumer 1 already knows $t$ and is doing so by acquiring information about what $j$ is. But consumer 2 is doing so without acquiring any information about what $j$ is or why it might be the right choice. What the state-action model accounts for is the fact that for each $t$, the probability of choosing $\xi_{j}=t$ differs from $1/2$ – what it would be if no information was acquired; what it doesn’t account for – or rather, treats as costless – is the information about $t$ itself. This issue can be formalized as follows.
Observational equivalence in state-action models when characteristics are state variables. In general, could an econometrician determine whether an insurance buyer in the state-action model has a priori private information about their future health expenditures $\theta$? Consider the canonical consumer problem setting (see e.g. [1]), the action space is $A=\{1,...,J\}$, and the utility vector is $\textbf{u}(\boldsymbol{\xi},\theta)=(u(1,\boldsymbol{\xi},\theta),...,u(J,\boldsymbol{\xi},\theta))$, which is a function of the random variables $j\in A$, $\boldsymbol{\xi}=(\xi_{1},...,\xi_{J})$, and $\theta\in\Theta$. The state-action model predicts that the optimal conditional choice probability is
$P(j|\boldsymbol{\xi},\theta)=P(j|\textbf{u}(\boldsymbol{\xi},\theta))=\frac{e^{u(j,\boldsymbol{\xi},\theta)}}{\sum_{k\in[J]}e^{u(k,\boldsymbol{\xi},\theta)}}$
This equation holds regardless of whether $\theta$ is assumed to be known to the consumer a priori or not. Thus, whether $\theta$ is known a priori is not identified in the state-action model.
Mathematically, we observe that by the law of iterated expectations,
$I_{P}(j,\textbf{u}):=\int\log\left(\frac{dP(j,\textbf{u})}{dP(j)dP(\textbf{u)}}\right)\>dP=\iint\log\left(\frac{dP(j,\textbf{u}|\theta)}{dP(j|\theta)dP(\textbf{u}|\theta)}\right)\>dP(j,\textbf{u}|\theta)\>dP(\theta)=\int I_{P}(j|_{\theta},\textbf{u}|_{\theta})\>dP(\theta)$
because $P(j|\theta)=P(j)$. So, the DM maximizes
$E_{P}[u]-\lambda I_{P}(j,\textbf{u})\equiv E_{P}\bigg{[}E_{P}[u|\theta]-\lambda I_{P}(j|_{\theta},\textbf{u}|_{\theta})\bigg{]}$
if and only if they maximize $E_{P}[u|\theta]-\lambda I_{P}(j|_{\theta},\textbf{u}|_{\theta})$ for almost every $\theta$. In words, the DM solves the state-action rational inattention problem overall if and only if they solve the state-action rational inattention problem in every health state $\theta$.
One final observation is that $I_{P}(j|_{\theta},\textbf{u}|_{\theta})=D_{KL}(P(\xi|\theta)\|\phi(\xi))$ for a natural prior $\phi$. To see this, suppose that out of the $J$ options, $k_{x}$ of them had the characteristic $x\in X$. A natural prior, then, would be $\phi(x)=k_{x}/J$. Without loss, suppose $\xi_{j}=x$. Then, $P(\xi_{j}|\theta,\boldsymbol{\xi})=P(x|\theta)/k_{x}$. Thus,
$\frac{P(j|\textbf{u})}{P(j)}=\frac{P(j|\theta,\boldsymbol{\xi})}{1/J}=\frac{P(\xi_{j}|\theta,\boldsymbol{\xi})}{1/J}=\frac{P(x|\theta)/k_{x}}{1/J}=\frac{P(x|\theta)}{\phi(x)}$
The need for consumer choice uncertainty. Given the non-identifiability problem, it may be tempting to simply assume that a consumer could observe the characteristics of their choices costlessly, and simply use the state-action model to model the uncertainty generated by $\theta$. Indeed, this is what most empirical discrete choice models implicitly do when $\xi$ is taken to be the choice variable. This is reasonable since the econometrician observes $P(\xi,\theta)$, not $P(k,\textbf{u})$. However, there are many circumstances in which the state-characteristic model we propose herein may be helpful.
For one, the choice probability predicted by the state-action model may be non-unique. Second, we may want to incorporate information that some options are more likely than others. Third, we may observe choices which appear strictly dominated, as is empirically common [2], which cannot be explained by modeling characteristics as actions in the state-action model. Fourth, there is no incorporation of the inherent ubiquity of certain options.
To illustrate these problems in structural econometric context, suppose Harold Zurcher is buying used buses, and $X=\{\text{red bus, blue bus, lemon bus}\}$ such that $\text{red}=\text{blue}\succ\text{lemon}$ in all states. In the state-action model without choice uncertainty, he would ever choose a lemon, no matter how abundant they were. And even if we knew red buses outnumbered blue buses 100-to-1, the state-action model does not prescribe a unique optimum policy – it could be optimal for Zurcher to only buy blue (or red).
So what are we missing? What we have shown is that when characteristics are modeled as state variables, the state-action imposes a cost function which is effectively $E_{P}[D_{KL}(P(\xi|\theta)\|\phi(\xi))]$, and when characteristics are modeled as actions, then naturally the cost function is $I_{P}(\xi,\theta)=E_{P}[D_{KL}(P(\theta|\xi)\|\mu(\theta))]$. In the former, the DM behaves as though observing $\theta$ is free. In the latter, the DM observes $\xi$ for free.
To wit, the natural extension of the state-action model which we propose in the next section contains two terms: $E_{P}[D_{KL}(P(\xi|\theta)\|\phi(\xi))]$ and $I_{P}(\xi,\theta)=E_{P}[D_{KL}(P(\theta|\xi)\|\mu(\theta))]$.
@@
Notation. Let $m,m^{\prime}$ be measures defined on the same measurable space. $m\ll m^{\prime}$ denotes absolute continuity. $m\ll m^{\prime}\ll m$ is interchangeable with $\sim m^{\prime}. m=m$$m^{\prime}$ if $m(A)=m^{\prime}(A)$ for each measurable set. The KL-divergence between two probabilities is defined
$D_{KL}(m\|m^{\prime})=\begin{cases}\int\log\left(\frac{dm}{dm^{\prime}}\right)\>dm&(m\ll m^{\prime})\\ \infty&\text{else}\end{cases}$
By convention, $\infty\cdot 0=0$ for integration purposes.222see [3]
@@
    (Bayes plausibility) $P|_{\theta}=\mu$
@@
    (marginal feasibility) $P|_{\xi}\ll\phi$
The cost function $\kappa_{\alpha,\lambda}$ can be written in several ways:
@@
Example: state-action with costlessly observed $\xi$. When $\alpha=0$, $\kappa_{\alpha,\lambda}(P)$ simplifies to $I_{P}(\xi,\theta)$, which corresponds to the state-action rational inattention model. Id est, when $\alpha=0$, the DM can choose characteristics directly, just as in the state-action model, the DM chooses actions directly.

### v2

@@ Abstract
We introduce a rational inattention model which produces a unique, interior, weighted multinomial logit conditional choice probability for an agent who acquires costly information about the hedonic characteristics (e.g. whether an insurance contract has high coverage) of their choices and about their payoff-relevant states (e.g. their risk of incurring a loss).
As usual, the objective is to choose a joint distribution subject to one marginal constraint (“Bayes plausibility”). We approach the problem by re-writing it in terms of an inner problem of maximizing over two constraints and an outer problem of choosing the “optimal constraint.” The inner problem is a Schrödinger bridge problem. The outer problem is strictly concave.
@@ Introduction
Many welfare-relevant choice problems feature uncertainty to both payoff-relevant states and menu characteristics. Consider insurance. A household may choose a contract that later appears ex-post “wrong”—too much or too little coverage. This could reflect uncertainty about the state (future medical expenditures), but it could also reflect limited attention to the menu (misunderstanding which plan is high coverage, misperceiving deductibles, or failing to screen out dominated options). These two failures have different normative implications: policies that improve forecasting of health shocks are not the same as policies that improve comprehension of contract characteristics.
This paper aims to build upon the state-action rational inattention model [matvejka2015rational], which has become a leading workhorse for disciplined departures from full information. In that framework, the decision-maker commits to a joint distribution over actions and states subject to Bayes plausibility, paying a cost linear in Shannon mutual information. The model yields a logit-like conditional choice rule and provides a clean mapping from information costs to stochastic choice. However, when the economic object of interest is choice among options with uncertain characteristics, the state–action approach confronts two limitations highlighted by the motivating insurance setting. First, it does not separate uncertainty about states from uncertainty about characteristics: treating characteristics as actions assumes they are costlessly observed; treating them as part of the state can render prior private information about other state components observationally irrelevant—a problem in decision problems with multi-dimensional uncertainty that we elaborate on in section˜2. Second, the state–action objective need not select a unique optimal marginal distribution over actions, and thus may fail to predict a unique conditional choice probability.
This paper proposes a generalized rational inattention model that directly targets these issues. The overarching idea is to model the econometrician’s observed joint distribution over characteristics and states as the result of a information policy which acquires information both (i) to match characteristics to states and (ii) to deviate from the baseline prevalence of characteristics in the environment. Formally, the decision-maker chooses an information policy $P$ over characteristics $\xi\in X$ and states $\theta\in\Theta$, subject to Bayes plausibility $P_{\theta}=\mu$ and feasibility $P_{\xi}\ll\phi$, where $\mu$ is the prior over states and $\phi$ is an exogenous prior over characteristics capturing their ubiquity in the market. The objective trades off expected utility against two information costs: the usual mutual information between $\xi$ and $\theta$ (learning about states via choice), and an additional KL divergence penalty that prices deviations of the marginal $P_{\xi}$ from $\phi$ (learning to screen the menu). A single parameter $\alpha\in[0,1]$ indexes the relative importance of these two margins of attention. The model nests the canonical state–action formulation at $\alpha=0$ and a “characteristics-attention” limit at $\alpha=1$.
This paper makes several contributions. First, it presents a tractable and testable stochastic-choice characterization. For intermediate $\alpha\in(0,1)$, optimal conditional choice probabilities satisfy a weighted multinomial logit:
$P(\xi|\theta)\propto\phi(\xi)^{\alpha}\,P_{\xi}(\xi)^{1-\alpha}\,e^{u(\xi,\theta)/\lambda}$ (1)
with an appropriate normalizing partition function. Equation (1) clarifies the economic forces behind stochastic choice: utility tilts choices toward high-payoff characteristics; $\phi$ captures the pull of prevalent characteristics (“lemons are harder to avoid when they are everywhere”); and the endogenous marginal $P_{\xi}$ captures the inertia induced by costly state learning (choices that are optimal in most states become focal unless attention is expended to condition sharply on $\theta$). The result parallels the celebrated logit characterization in [matvejka2015rational] while adding a new, economically meaningful weighting by $\phi$.
Second, it presents a rational inattention model that yields a point prediction. In contrast to the state-action model, a major advantage of this model for $\alpha\in(0,1)$ is that it delivers a unique optimal marginal $P_{\xi}$ and hence a unique conditional choice rule. Moreover, the solution is interior relative to $\phi$. This feature both improves empirical discipline by providing point predictions and rationalizes phenomena that are difficult to accommodate when characteristics are assumed perfectly observed, such as selection of dominated or low-quality options when such options are common.
Finally, this paper generalizes rational inattention to information policies over non-discrete spaces and presents a novel connection between rational inattention and recent developments in entropic optimal transport. Fixing the marginal $P_{\xi}=\nu$ converts the inner problem into an entropic optimal-transport problem (a Schrödinger bridge) between $\nu$ and $\mu$. Because the Schrödinger bridge is unique for every $\nu$, the problem simplifies to finding the optimal $\nu$. This turns out to be a strictly convex optimization problem.
Section 2 motivates the model by formalizing why the standard state–action approach cannot simultaneously discipline both observed information gain in $P(\xi|\theta)$ relative to a characteristic baseline and observed informativeness of choices about states. Section 3 introduces the state–characteristic model on general Polish spaces, defines the two-part information cost, and discusses the nesting cases $\alpha=0$ and $\alpha=1$. Section 4 states the main characterization results: the Gibbs property and the weighted logit formula, along with uniqueness and interiority. Section 5 derives testable restrictions and the overidentification logic using entry-driven shifts in $\phi$. Sections 68 develop the variational and optimal-transport machinery: first-step orthogonality, the Schrödinger bridge formulation of the constrained problem, and the strictly concave marginal program that delivers existence and uniqueness. We conclude with a section 10 which discusses computation in finite spaces.
Notation. We use $X$ to denote the space of characteristics and $\Theta$ to denote the space of states. $\xi$ is a generic element of $X$, typically representing the characteristic of a choice of the decision-maker. $\boldsymbol{\xi}=(\xi_{1},...,\xi_{J})$ is a random vector that represents the characteristics of a menu $A=\{1,...,J\}$. $P$ denotes an information policy, typically the optimal one. $E_{Q}$ is the expectation with respect to a measure $Q$. For a joint distribution $Q$ over random elements $Y$ and $Z$, let $Q_{Y}$ and $Q_{Z}$ denote the respective marginals. $Q\ll R$ means that $Q$ is absolutely continuous with respect to $R$. $D_{KL}(Q\|R)=E_{Q}\left[\log\left(\frac{dQ}{dR}\right)\right]$ is the information gain of $Q$ over $R$, commonly known as the Kullback-Leibler divergence. $I_{Q}(Y,Z)=D_{KL}(Q_{Y,Z}\|Q_{Y}\otimes Q_{Z})$ is the mutual (Shannon) information between random elements $Y\in S$ and $Z\in T$ under joint distribution $Q$. $\Delta S$ represents the set of all probability distributions over $S$. $\Pi(\cdots)$ is the set of all joint distributions satsifying the marginal restrictions $(\cdots)$.
@@
In short, the problem with modeling inattention to states and characteristics using the state-action model boils down to the fact that the Shannon information between actions $a$ and random elements $(\omega_{1},\omega_{2})$ does not account for the information between $(a,\omega_{1})$ and $\omega_{2}$. In the context of consumer choice over products with arbitrary indices $j\in\{1,...,J\}$, characteristics $\boldsymbol{\xi}=(\xi_{1},...,\xi_{J})$, and states $\theta$, this is to say the Shannon information between $j$ and $(\boldsymbol{\xi},\theta)$ does not account for information between $\theta$ and $(j,\boldsymbol{\xi})$.
As a simple example, consider a canonical state-action consumer choice problem. We have a health insurance market with two insurance contracts $j\in\{1,2\}$. A consumer faces two sources of uncertainty. First, he will be high risk ($\theta=HR$) with probability (w.p.) $\frac{1}{2}$ and low risk ($\theta=LR$) w.p. $\frac{1}{2}$. Second, he knows that one contract has high coverage and the other has low coverage but is uncertain as to how they are permuted: $\boldsymbol{\xi}\equiv(\xi_{1},\xi_{2})=(HC,LC)$ w.p. $\frac{1}{2}$ and $\boldsymbol{\xi}\equiv(\xi_{1},\xi_{2})=(LC,HC)$ w.p. $\frac{1}{2}$. The consumer receives utility 1 if he chooses “correctly” ($HC$ when $\theta=HR$ and $LC$ when $\theta=LR$) and 0 otherwise.
In the state-action model, the consumer’s information cost from committing to an information policy $Q(j,(\boldsymbol{\xi},\theta))$ is given by the mutual information $I_{Q}$ between the action $j$ and the random element $(\boldsymbol{\xi},\theta)$. The cost of an information policy $Q$ which sends the “correct” action recommendation signal with probability $q>\frac{1}{2}$ and the “incorrect” signal with probability 1q1-q is
$\kappa(Q)=I_{Q}(j,(\boldsymbol{\xi},\theta))=\log 2+q\log q+(1-q)\log(1-q)$ (2)
where $I_{Q}(\cdot,\cdot)$ is mutual information under probability measure $Q$.
The problem is that a consumer who already knows his $\theta$ and is only uncertain about $\boldsymbol{\xi}$ also has an optimal information policy that involves receiving the “correct” signal with probability $q$. The cost of such a policy is also given by eq.˜2. Because the consumer who knows $\theta$ and the consumer who does not both face the exact same objective function, they end up being observationally identical.
For the consumer who knows $\theta$ already, the state-action information cost function correctly penalizes the information that allows the consumer to choose the “correct” contract with better than random odds – that is, the information gain of $P(\xi|\theta)$ over $\phi(\xi)$, where $\xi$ represents the chosen contract and $\phi(\xi)=\frac{1}{2}$ is the probability of choosing $\xi$ randomly.
What the information cost $I_{Q}(j,(\boldsymbol{\xi},\theta))$ fails to account for is the information gain in $\theta$. This is because when the consumer who does not know $\theta$ receives the signal “choose option j=1j=1” from the oracle, he receives no per se information about his future health risk222in this case, he also receives no information about the contracts – and hence does not have to pay for such information under the state-action cost function.
The key insight is that although the action recommendation signal $j$ does not provide any information $\theta$ (and $\boldsymbol{\xi}$ does not either), $(j,\boldsymbol{\xi})$ does. In our health insurance example, “choose option $j=1$ where $\xi_{1}$ is likely to be HCHCdoes indeed provide information that the state is likely $HR$. That is, what we are missing in the state-action cost function is $I_{Q}(\theta,(j,\boldsymbol{\xi}))$ – the mutual information between $\theta$ and $(j,\boldsymbol{\xi})$. As it turns out, under the optimal information policy, it suffices to account for the mutual information between $\theta$ and the characteristics $\xi$ of the chosen contract. To wit, we present a model which accounts for this.
@@
We now present a model of rational inattention in which information about states and characteristics are both costly. While our model is microfounded by a decision problem with discrete actions $j\in A=\{1,...,J\}$ and random elements $(\xi_{1},...,\xi_{J},\theta)\in X^{J}\times\Theta$, the utility of the decision-maker depends only on the state $\theta$ and the chosen characteristic $\xi\equiv\xi_{j}$. Thus, we can choose to either model the decision-maker’s information policy over $A\times X^{J}\times\Theta$ or, more directly, over $X\times\Theta$. We go with the latter for several reasons. The foremost reason is that it is more tractable theoretically. But it is also more tractable in applied settings, as econometricians typically have data on the state $\theta$ and the chosen characteristic $\xi$, but not necessarily on $(j,\boldsymbol{\xi},\theta)$. Finally, this formulation has the added bonus of being interpretable as an interpolation between two state-action models: one in which $\xi$ is costless to learn, and one in which $\theta$ is costless to learn. Despite the tractability, the model is nevertheless abstract: we let the characteristic space and the state space be general Polish spaces.
@@
    (Bayes plausibility) $P_{\theta}=\mu$
@@
    (marginal feasibility) $P_{\xi}\ll\phi$
The cost function $\kappa_{\alpha,\lambda}$ is defined
[…]
The cost function can also be written equivalently in several equivalent ways:
@@
Microfoundation. Consider a decision problem defined by actions $j\in A=\{1,...,J\}$, random elements $(\boldsymbol{\xi},\theta)\in X^{J}\times\Theta$, an exogenous prior $\rho\in\Delta(X^{J}\times\Theta)$, and a utility function $u(\xi,\theta)$ that depends only on the state $\theta$ and the chosen characteristic $\xi$. Assume also that:
  1. 1.
    (independence) $\boldsymbol{\xi}\bot\!\!\!\bot\theta$
  2. 2.
    (a priori homogeneity) the indexing of choices $\{1,...,J\}$ are arbitrary
  3. 3.
    (design-based uncertainty) the DM knows the number of choices $\phi(x)$ that correspond to characteristic $x\in X$, but does not know how these characteristics are permuted among the options 1,,J1,...,J
The DM’s objective is to choose an information policy $P\in\Delta(A\times X^{J}\times\Theta)$ which maximizes expected utility less an information cost subject to Bayes plausibility $P_{X^{J}\times\Theta}=\rho$.
Where this decision problem departs from the state-action model is that the DM in this decision problem faces an information cost with two terms. First, there is the (usual) Shannon information between actions $j$ and random elements $(\boldsymbol{\xi},\theta)$. We can re-write this as
$I_{P}(j,(\boldsymbol{\xi},\theta))$ $=E_{P}\left[\log\left(\frac{P(j,(\boldsymbol{\xi},\theta))}{P(j)\times P((\boldsymbol{\xi},\theta))}\right)\right]$
$=E_{P}\left[\log\left(\frac{P(j,(\boldsymbol{\xi},\theta)|\theta)}{P(j)\times P(\boldsymbol{\xi}|\theta)}\right)\right]$
$=E_{P}\left[\log\left(\frac{P(j,(\boldsymbol{\xi},\theta)|\theta)}{P(j|\theta)\times P(\boldsymbol{\xi}|\theta)}\right)\right]$ (homogeneity)
$=E_{P}\left[\log\left(\frac{P(\boldsymbol{\xi}|j,\theta)}{P(\boldsymbol{\xi}|\theta)}\right)\right]$
$=E_{P}\left[\log\left(\frac{P(\boldsymbol{\xi}|j,\theta)}{P(\boldsymbol{\xi})}\right)\right]$ (independence)
Let $\xi_{-j}=(\xi_{1},...,\xi_{j-1},\xi_{j+1},...,\xi_{J})$. By complementarity under the permutation assumption, $\xi_{j}$ pins down what characteristics are left in $\xi_{-j}$, but does not pin down how they are permuted. If the information policy is optimal, $j$ should provide no additional information about how they are permuted as it is not payoff-relevant by assumption. Thus, if $P$ is optimal, then for any $\textbf{x}=(x_{1},...,x_{J})\in X$,
$\frac{P(\boldsymbol{\xi}=\textbf{x}|j,\theta)}{P(\boldsymbol{\xi}=\textbf{x})}\equiv\frac{P(\boldsymbol{\xi}=(x_{j},x_{-j})|j,\theta)}{P(\boldsymbol{\xi}=(x_{j},x_{-j}))}=\frac{P(\xi=x_{j}|j,\theta)P(\xi_{-j}=x_{-j}|\xi_{j},j,\theta)}{P(\xi_{j}=x_{j})P(\xi_{-j}=x_{-j}|\xi_{j})}=\frac{P(\xi_{j}=x_{j}|j,\theta)}{P(\xi_{j}=x_{j})}\equiv\frac{P(\xi=x_{j}|\theta)}{\phi(x_{j})}$
where $\xi$ is the chosen characteristic. Hence the first term of the DM’s information cost can be re-written as
$I_{P}(j,(\boldsymbol{\xi},\theta))=E_{P}\left[\log\left(\frac{P(\xi|\theta)}{\phi(\xi)}\right)\right]=E_{P}[D_{KL}(P(\xi|\theta)\|\phi(\xi))]$
The second information cost faced by the DM is for the expected information gain from $P(\theta|j,\boldsymbol{\xi})$ over the exogenous prior $P(\theta)$. Again, if the information policy is optimal, then $P(\theta|j,\boldsymbol{\xi})=P(\theta|\xi_{j})$; that is, the chosen $\xi_{j}$ should provide information about $\theta$, but the exact permutation of $\xi_{-j}$ should not, since it is not payoff relevant. Hence, this second cost term can be re-written as
$I_{P}(\theta,(j,\boldsymbol{\xi}))E_{P}[D_{KL}(P(\theta|j,\boldsymbol{\xi})\|P(\theta))]=E_{P}[D_{KL}(P(\theta|\xi)\|P(\theta))]$
Setting the DM’s cost function to be a weighted sum of these two cost terms yields
$\kappa_{\alpha,\lambda}(P)$ $=\alpha\lambda I_{P}(j,(\boldsymbol{\xi},\theta))+(1-\alpha)\lambda I_{P}(\theta,(j,\boldsymbol{\xi}))$
$=\alpha\lambda E_{P}[D_{KL}(P(\xi|\theta)\|\phi(\xi))]+(1-\alpha)\lambda E_{P}[D_{KL}(P(\theta|\xi)\|P(\theta))]$
which is precisely the cost function presented in our model.
@@
Example: state-action with costlessly observed $\boldsymbol{\xi}$. When $\alpha=0$, $\kappa_{\alpha,\lambda}(P)$ simplifies to $I_{P}(\xi,\theta)$, which is exactly the cost function we would get if we treated characteristics as costlessly observed and then applied the state-action model.